During a hurricane, the atmospheric pressure inside a house may blow off the roof because of the reduced pressure outside. If air (density 1.293 kg/m3) is blowing over the roof of a house at a speed of 142 km/hr, what is the pressure difference between the inside and the outside of the house?
Answer #1

Solution)

Given,

Speed, v1=142 km/hr= 39.47 m/s

Speed, v2=0

Density, rho=1.293 kg/m^3

According to Bernoullie's Theorem,

Pressure difference=0.5*Rho*(v1^2- v2^2)

Delta P=0.5*(1.293)*(39.47)^2=1007.47 Pa (Ans)

============